# Correlation between the charge radii difference in mirror partner nuclei and the symmetry energy slope

### v1

@@ II THEORETICAL FRAMEWORK
For further focusing on the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$, charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar against slope parameter $L$ are shown in Fig. 2. In this figure, clear linear correlations can be observed for these mirror-pair nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar. For the volume-type pairing force, the Pearson coefficients are almost no change in the mirror partner nuclei 32Ar-32Si, 36Ca-36S and 38Ca-38Ar, namely the corresponding $R$ values are 0.926, 0.881 and 0.836, respectively. With considering the mixed-type pairing force, the Pearson coefficients $R$ are slightly improved with respect to the no pairing case. Actually, the Pearson coefficients $R$ are rather poor if the surface pairing force is taken into account. Here the corresponding figures are not shown, but the Pearson coefficients $R$ with considering surface-type pairing force are $0.887$, $0.791$ and $0.743$ for the corresponding mirror partner nuclei 32Ar-32Si, 36Ca-36S and 38Ca-38Ar, respectively. This means that pairing effect can decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ if the pairing interactions are only tackled by the surface-type pairing force.
@@
Figure 3: (Color online) Plots for the values of slope parameter $L$ extracted in this work (54Ni-54Fe, 38Ca-38Ar, 36Ca-36S, and 32Ar-32Si) and those from the existing literatures. The partially enumerated values of $L$, such as Wang $et~{}al$Wang and Li (2013), Brown $et~{}al$Brown et al. (2020), Pineda $et~{}al$Pineda et al. (2021), An $et~{}al$An et al. (2023c), Bano $et~{}al$Bano et al. (2023), König $et~{}al$König et al. (2023), pygmy dipole resonance (PDR) Carbone et al. (2010), electric dipole polarizability $\alpha_{\mathrm{D}}$ Roca-Maza et al. (2015), GW170817 Raithel and Özel (2019), pion production ($\pi^{-}/\pi^{+}$iu et al. (2023), and the charge exchange spin-dipole (SD) excitations Cheng et al. (2023), are also given. The color band represents the extracted slope parameter $L=42.57$-$50.64$ MeV in this literature.
@@
As mentioned in Refs. Wang and Li (2013); Brown (2017); Brown et al. (2020); Pineda et al. (2021); König et al. (2023), charge radii differences in mirror partner nuclei provide an alternative approach to pin down the isovector components in the equation of state of asymmetric nuclear matter. To facilitate the relatively clear linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$, the mixed-type pairing force is employed in the following discussion. In Fig. 3, the deduced slope parameters $L$ from the differences of charge radii of mirror-pair nuclei 54Ni-54Fe, 38Ca-38Ar, 36Ca-36S, and 32Ar-32Si are drawn by shaping the covered range in Fig. 2. Combing the experimental data, the extracted slope parameter falls into the range of $L=42.57$-$50.64$ MeV. From this figure, one can find that the deduced value of $L$ covers the uncertainty range deriving from the mirror charge radii differences Wang and Li (2013); Brown et al. (2020); Pineda et al. (2021); An et al. (2023c); Bano et al. (2023); König et al. (2023). Here it is mentioned that the upper limit range of $L$ coming from 32Ar-32Si is 96.65 MeV. This result is overestimated in comparison with the Ref. König et al. (2023) where the upper range is confined to be $L\leq 60$ MeV. This may be due to the fact that shape deformation is ignored in our study. Besides, the pairing correlations and the limited Skyrme parameter sets used here may also have an influence on evaluating the interval range of slope parameter $L$.
@@
Also, the value of symmetry energy slope $L$ extracted from pygmy dipole resonance (PDR) locates at the internal range of $L=64.8\pm 15.7$ Carbone et al. (2010). The covered range is very narrow with respect to other cases of this figure. Furthermore, the extracted $L$ in this work can cover the uncertainty ranges deriving from the nuclear electric dipole polarizability $\alpha_{\mathrm{D}}$ Roca-Maza et al. (2015), dense astrophysical event GW170817 Raithel and Özel (2019), the pion production ratios in heavy-ion collisions Liu et al. (2023) and the charge exchange spin-dipole (SD) excitations Cheng et al. (2023). Besides, this deduced result can mostly cover the range of $L$ shown in Ref. Zhang and Chen (2014), in which the slope parameter captures the range of $L=47.3\pm 7.8$. In a word, the extracted range of slope parameter $L=42.57$-$50.64$ MeV means that the rather soft equation of state of isospin asymmetric nuclear matter can be obtained from the implication of charge radii difference in mirror-pair nuclei.
@@
Figure 4: (Color online) Charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 36Ca-36S and 38Ca-38Ar are plotted as a function of slope parameter $L$ classified by various isoscalar incompressibility coefficients $K$. The corresponding experimental data are taken from Ref. Brown et al. (2020).
@@
As shown in Fig. 2, the Pearson coefficients become lower for the mirror-pair nuclei 36Ca-36S and 38Ca-38Ar. This suggests that the profoundly theoretical uncertainties are encountered in the calibrating protocol. In order to clarify this phenomenon, the charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 36Ca-36S and 38Ca-38Ar against the slope parameter $L$ classified by various isoscalar incompressibility coefficients $K$ are shown in Fig. 4. From both panels of Fig. 4, one can find that the isoscalar incompressibility has an influence on the charge radii difference of mirror partner nuclei. Even from $K\approx 230$ to $K\approx 250$ MeV, the deviated range for $L$ is about $38.0$ MeV. Here, as shown in Fig. 3, we mention that the influence of the isoscalar incompressibility coefficient cannot be taken into account in evaluating the slope parameter $L$. Significantly the isoscalar incompressibility may be indispensable in calibrating protocol. As is well known, the accurate description of the density-dependent behavior of symmetry energy plays an important role in recognizing various physical mechanisms through nuclear models, such as the production of superheavy elements Veselsky et al. (2016), and the cluster radioactivity Xu et al. (2014); Yang et al. (2023b), etc. Especially, recent study suggests that the effective proton-neutron chemical potential difference of neutron-rich nuclei is strongly sensitive to the symmetry energy Qiu et al. (2023). Therefore, more aspects should be taken into account in constraining the slope parameter $L$.
@@
In this work, the correlation between the charge radii differences ($\Delta{R_{\mathrm{ch}}}$) in mirror-pair nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe and slope parameter $L$ at saturation density is investigated by the spherical Skyrme energy density functionals. The calculated results suggest that the surface pairing force can further decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$. The deduced slope parameter locates the interval range of $L=42.57$-$50.64$ MeV under the mixed pairing interaction, namely the rather soft equation of state. This is in accord with most existing literatures. Meanwhile, the influence of isoscalar nuclear matter property on $\Delta{R_{\mathrm{ch}}}$ is inevitable in constraining the nuclear matter EoS. This may suggest that most parametrization forces are ruled out in the calibrated procedure. As shown in Fig. 4, the correlation between charge radii difference of mirror nuclei and the slope parameter of symmetry energy is influenced by nuclear incompressibility. This seems to suggest that the Skyrme parameter sets adjusted to different slope parameters under specific incompressibility coefficient can give highly linear correlation between the charge radii difference of mirror partner nuclei and the slope parameter $L$. Meanwhile, this may provide an alternative approach to evaluate the possible influential quantities in determining the EoS of asymmetric nuclear matter. Shown in Ref. Huang et al. (2023), the calculated results obtained by density energy functionals consisting of Skyrme and covariant models suggest that more mirror-pair nuclei are ruled out in constraining the slope parameter $L$, but the superior candidates of mirror-pair nuclei 44Cr-44Ca and 46Fe-46Ca may be used to ascertain the limit range of $L$. Thus more discussion should be performed in the proceeding work.
Recently studies show that the charge-changing cross section measurements can be used to determine the EoS of nuclear matter with mirror nuclei Xu et al. (2022b). This method can also be employed to investigate the charge radii of exotic nuclei Wang et al. (2023) and the isospin interactions Zhao et al. (2023). This provides an accessable approach to ascertain the nuclear symmetry energy. From the aspect of the charge radii difference in mirror-pair nuclei, the underlying mechanism should be taken into account in determining the nuclear symmetry energy, such as the local relations among the adjacent neighbouring nuclei Sun et al. (2014); Bao et al. (2016, 2020); Ma et al. (2021), the $\alpha$-cluster formation Qian et al. (2014); Ni et al. (2013), etc. Thus more high-precision charge radii data are urgently required in experimental and theoretical researches.

### v2

@@ II THEORETICAL FRAMEWORK
For further focusing on the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$, the charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar against slope parameter $L$ are shown in Fig. 2. In this figure, clear linear correlations can be observed for these mirror-pair nuclei (32Ar-32Si, 36Ca-36S, and 38Ca-38Ar). For the volume-type pairing force, the Pearson coefficients show almost no change in the mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar, with $R$ values of 0.926, 0.881, and 0.836, respectively. Considering the mixed-type pairing force, the Pearson coefficients $R$ are slightly improved with respect to the no pairing case. Actually, the Pearson coefficients $R$ are rather poor if the surface pairing force is taken into account. The corresponding figures are not shown here, but the Pearson coefficients $R$ for the surface-type pairing force are $0.887$, $0.791$, and $0.743$ for the mirror partner nuclei 32Ar-32Si, 36Ca-36S, and 38Ca-38Ar, respectively. This means that the pairing effect can decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$ if the pairing interactions are only addressed by the surface-type pairing force.
@@
Figure 3: (Color online) Plots of the values of slope parameter $L$ obtained in this work (54Ni-54Fe, 38Ca-38Ar, 36Ca-36S, and 32Ar-32Si) and in previous studies. The partially enumerated values of $L$ from Wang $et~{}al$Wang and Li (2013), Brown $et~{}al$Brown et al. (2020), Pineda $et~{}al$Pineda et al. (2021), An $et~{}al$An et al. (2023c), Bano $et~{}al$Bano et al. (2023), König $et~{}al$König et al. (2024), pygmy dipole resonance (PDR) Carbone et al. (2010), electric dipole polarizability $\alpha_{\mathrm{D}}$ Roca-Maza et al. (2015), GW170817 Raithel and Özel (2019), pion production ($\pi^{-}/\pi^{+}$iu et al. (2023), and charge exchange spin-dipole (SD) excitations Cheng et al. (2023), are also presented. The color band represents the slope parameter $L=42.57$-$50.64$ MeV obtained in this study.
@@
As mentioned in Refs. Wang and Li (2013); Brown (2017); Brown et al. (2020); Pineda et al. (2021); König et al. (2024), charge radii differences in mirror partner nuclei provide an alternative approach to pin down the isovector components in the EoS of asymmetric nuclear matter. To analyze the relatively clear linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$, the mixed-type pairing force is employed in the following discussion. In Fig. 3, the deduced slope parameters $L$ from the differences of charge radii of mirror-pair nuclei 54Ni-54Fe, 38Ca-38Ar, 36Ca-36S, and 32Ar-32Si are drawn by shaping the range covered in Fig. 2. Combing the experimental data, the extracted slope parameter falls into the range of $L=42.57$-$50.64$ MeV. From this figure, one can find that the deduced value of $L$ covers the uncertainty range derived from the mirror charge radii differences Wang and Li (2013); Brown et al. (2020); Pineda et al. (2021); An et al. (2023c); Bano et al. (2023); König et al. (2024). Here, the upper limit range of $L$ coming from 32Ar-32Si is 96.65 MeV. This result is overestimated in comparison with that in Ref. König et al. (2024), where the upper range is confined to $L\leq 60$ MeV. This may be due to the fact that shape deformation is ignored in our study. Besides, the pairing correlations and the limited Skyrme parameter sets used here may also have an influence on evaluating the interval range of slope parameter $L$.
@@
The value of the symmetry energy slope $L$ extracted from the pygmy dipole resonance (PDR) locates at the interval range of $L=64.8\pm 15.7$ MeV Carbone et al. (2010). The covered range is very narrow compared with that of the other cases depicted in this figure. Furthermore, the $L$ obtained in this work can cover the uncertainty ranges derived from the nuclear electric dipole polarizability $\alpha_{\mathrm{D}}$ Roca-Maza et al. (2015), dense astrophysical event GW170817 Raithel and Özel (2019), pion production ratios in heavy-ion collisions Liu et al. (2023), and charge exchange spin-dipole (SD) excitations Cheng et al. (2023). Besides, this deduced result covers most of the range of $L$ presented in Ref. Zhang and Chen (2014), in which the slope parameter captures the range of $L=47.3\pm 7.8$ MeV. In summary, the extracted range of slope parameter $L=42.57$-$50.64$ MeV means that the rather soft EoS of the isospin asymmetric nuclear matter can be obtained from the implication of charge radii difference in mirror-pair nuclei.
@@
Figure 4: (Color online) Charge radii differences $\Delta{R_{\mathrm{ch}}}$ of mirror partner nuclei 36Ca-36S and 38Ca-38Ar plotted as a function of slope parameter $L$ classified by various isoscalar incompressibility coefficients $K$. The corresponding experimental data are taken from Ref. Brown et al. (2020).
@@
As shown in Fig. 2, the Pearson coefficients become lower for the mirror-pair nuclei 36Ca-36S and 38Ca-38Ar. This suggests that the profoundly theoretical uncertainties are encountered in the calibrating protocol. To clarify this phenomenon, the charge radii differences $\Delta{R_{\mathrm{ch}}}$ of the mirror partner nuclei 36Ca-36S and 38Ca-38Ar against the slope parameter $L$ classified by various isoscalar incompressibility coefficients $K$ are shown in Fig. 4. From both panels of Fig. 4, one can find that the isoscalar incompressibility has an influence on the charge radii difference of mirror partner nuclei. Even from $K\approx 230$ to $K\approx 250$ MeV, the deviated range for $L$ is approximately $38.0$ MeV. Here, as shown in Fig. 3, we mention that the influence of the isoscalar incompressibility coefficient cannot be taken into account in evaluating the slope parameter $L$. Significantly the isoscalar incompressibility may be indispensable in the calibrating protocol. As is well known, the accurate description of the density-dependent behavior of symmetry energy plays an important role in recognizing various physical mechanisms through nuclear models, such as the production of superheavy elements Veselsky et al. (2016) and cluster radioactivity Xu et al. (2014); Yang et al. (2023b). In particular, a recent study suggests that the effective proton-neutron chemical potential difference of neutron-rich nuclei is strongly sensitive to the symmetry energy Qiu et al. (2024). Therefore, more aspects should be taken into account in constraining the slope parameter $L$.
@@
In this work, the correlation between the charge radii differences ($\Delta{R_{\mathrm{ch}}}$) in mirror-pair nuclei 32Ar-32Si, 36Ca-36S, 38Ca-38Ar, and 54Ni-54Fe and slope parameter $L$ at saturation density is investigated based on the spherical Skyrme EDFs. The calculated results suggest that the surface pairing force can further decrease the linear correlation between $\Delta{R_{\mathrm{ch}}}$ and $L$. The deduced slope parameter locates at the interval range of $L=42.57$-$50.64$ MeV under the mixed pairing interaction, that is, the rather EoS. This is in aggrement with most of the previous studies. Meanwhile, the influence of the isoscalar nuclear matter property on $\Delta{R_{\mathrm{ch}}}$ inevitably constrains the nuclear matter EoS. This may suggest that most of the parametrization forces are ruled out in the calibrated procedure. As shown in Fig. 4, the correlation between the charge radii difference of mirror nuclei and the slope parameter of symmetry energy is influenced by nuclear incompressibility. This seems to suggest that the Skyrme parameter sets adjusted to different slope parameters under a specific incompressibility coefficient can result in a highly linear correlation between the charge radii difference of mirror partner nuclei and the slope parameter $L$. Meanwhile, this may provide an alternative approach to evaluate the possible influential quantities in determining the EoS of asymmetric nuclear matter. As shown in Ref. Huang et al. (2023), the calculated results obtained by density energy functionals consisting of Skyrme and covariant models suggest that more mirror-pair nuclei are ruled out in constraining the slope parameter $L$, but the superior candidates of mirror-pair nuclei 44Cr-44Ca and 46Fe-46Ca may be used to ascertain the limit range of $L$. Thus, more discussion should be performed in the proceeding work.
Recent studies have shown that the charge-changing cross section measurements can be used to determine the EoS of nuclear matter with mirror nuclei Xu et al. (2022b). This method can also be employed to investigate the charge radii of exotic nuclei Wang et al. (2023) and the isospin interactions Zhao et al. (2023). This provides an accessable approach to ascertain the nuclear symmetry energy. From the aspect of the charge radii difference in mirror-pair nuclei, the underlying mechanism should be taken into account in determining the nuclear symmetry energy, such as the local relations among the adjacent neighbouring nuclei Sun et al. (2014); Bao et al. (2016, 2020); Ma et al. (2021) and $\alpha$-cluster formation Qian et al. (2014); Ni et al. (2013). Thus more high-precision charge radii data are urgently required in experimental and theoretical studies.